OpenStudy (anonymous):

The molar mass of a metal (M) is 65.4g/mol and it forms a chloride of unknown composition. Electrolysis of a sample of the molten chloride with a current of 6.42A for 23.6 minutes produces 3.08g of M at the cathode. Find the empirical formula of the chloride.

OpenStudy (anonymous):

the mass that generated on cathod is m=C*A*I*t/n you need to know the n this is valence C is constant A is atomic mass so n= C*A*I*t/m=2 your formula of cloridde : \[MeCl _{2}\]